What is the slope of the line perpendicular to 4x + 5 = 3y – 6 +x?

A. 1
B. -1/4
C. -1
D. -1/5

Answer by Alan3354(69443)

4x + 5 = 3y – 6 +x?
Only the x and y terms are of interest:
4x = 3y-x
3x = 3y
slope = 3/3 = 1
The slope of lines perpendicular (not just "the line") is the negative inverse,
= -1
-----------------
The coefficients 5 and -6 have no effect on the slope.